If the average price of a new one-family home is $246,300 with a standard deviation of $15,000, find the minimum and maximum prices of the houses that a contractor will build to satisfy the middle 80% of the market.Assume the variable is normally distributed.
Answer by stanbon(75887)

If the average price of a new one-family home is $246,300 with a standard deviation of $15,000, find the minimum and maximum prices of the houses that a contractor will build to satisfy the middle 80% of the market. Assume the variable is normally distributed.
------------------------------------
Find the z-values that bracket 80% around the mean.
------
Left tail of 10% = -1.2816
Right tail of 10% = +1.2816
------------------------------
minimum price = -1.2816*15000+ 246,300 = $227,076.73
maximum price = +1.2816*15000 + 246,300 = $265,524.00
======================
Cheers,
Stan H.